### v2

# Anatomy of a Digital Bubble: Lessons Learned from the NFT and Metaverse Frenzy

Abstract
In the past few years, “metaverse” and “non-fungible tokens (NFT)” have become buzzwords, and the prices of related assets have exhibited large fluctuations. Are those characteristic of a speculative bubble? In this paper, we attempt to answer this question, and better understand the underlying economic dynamics. Specifically, we look at Decentraland, a virtual world platform where land parcels are sold as NFT collections. We find that initially, land prices followed traditional real estate pricing models – in particular, value decreased with distance from the most desirable areas – suggesting Decentraland behaved much like a virtual city. However, these real estate pricing models stopped applying when both the metaverse and NFTs gained increased popular attention and enthusiasm in 2021, suggesting a new driving force for the underlying asset prices. At that time, following a substantial rise in NFT market values, short-term holders of multiple parcels began to take major selling positions in the Decentraland market, which hints that, rather than building a metaverse community, early Decentraland investors preferred to cash out when land valuations became inflated. Our analysis also shows that while the majority of buyers are new entrants to the market (many of whom joined during the bubble), liquidity (i.e., parcels) was mostly provided by early adopters selling, which caused stark differences in monetary gains. Early adopters made money – more than 10,000 USD on average per parcel sold – but users who joined later typically made no profit or even incurred losses in the order of 1,000 USD per parcel. Unlike established markets such as financial and real estate markets, newly emergent digital marketplaces are mostly self-regulated. As a result, the significant financial risks we identify indicate a strong need for establishing appropriate standards of business conduct and improving user awareness.
1 Introduction
The word “metaverse” was originally coined in science fiction in 1992 [36], to denote an extension of the Internet into a single virtual world. With advances in computer graphics and virtual reality hardware, we have seen increasingly realistic attempts to build a metaverse – or at least a number of virtual worlds – over the past couple of decades, with platforms such as Second Life becoming quite popular. Tellingly, in a bid that virtual worlds were about to harness enormous economic potential, in October 2021, Facebook rebranded itself as “Meta,” hoping to become one of the main building blocks of this metaverse [45].
Until recently, most virtual worlds (e.g., Second Life, World of Warcraft, etc.) were centralized, and operated by a single company. Although second-hand markets exist, in traditional, centralized virtual worlds, users willing to purchase wearables for instance to decorate their virtual-world avatars or acquire virtual land, usually have to buy them from the hosting company.
On the other hand, newer projects, such as Decentraland [^1] https://decentraland.org/. or The Sandbox, [^2] https://www.sandbox.game/en/. have leveraged recent advances in decentralization technologies, e.g., “blockchains” and cryptocurrencies that instantiate bankless peer-to-peer payment systems. These newer projects rely on a blockchain to host vital records of their underlying virtual worlds. Further, governance, instead of being centralized, is handled by decentralized anonymous organizations (DAOs). Roughly speaking, DAOs attempt to implement a democratic process where governance tokens can be purchased by the general public. Similar to holding company shares, votes are typically weighted by the amount of governance tokens one holds [37]. Likewise, decentralized virtual world items (e.g., wearables, land, and so forth) are often instantiated as “non-fungible tokens” (NFTs), which are also supported by (distributed) blockchain infrastructure rather than a hosting company. Importantly, both governance tokens and NFTs can be traded outside the virtual world platform, e.g., on cryptocurrency exchanges, or third-party NFT marketplaces such as OpenSea. [^3] https://opensea.io.
Decentralized virtual worlds appear to have capitalized on the surge in popularity of both the metaverse and NFTs. Concretely, in late 2021, the renewed interest in the metaverse following Facebook’s announcement, combined with the simultaneous sudden and substantial rise in cryptocurrency prices, caused a massive increase in Decentraland virtual land (parcel) prices, and in the valuation of its governance token called MANA. The median sales price of land in Decentraland rose from less than 1,000 USD in 2020 to roughly 15,000 USD in Q4 of 2021—only to plummet back to 5,000 USD in the subsequent half year. By the end of 2023, prices had gone back down to approximately 1,000 USD, and, as we will see, activity was minimal. At the time of writing, parcels go for approximately 200 USD.
Such volatility shows the hallmarks of an economic bubble. Was that indeed the case? How did the situation affect the wealth of all investors participating in Decentraland—both buyers and sellers? Were certain groups of users disproportionately and/or adversely affected? If so, could regulation help “level the playing field?” Our main contribution is addressing these questions based on a quantitative analysis of online conversations on Reddit and of Ethereum blockchain records about Decentraland.
We argue this is an important contribution to the study of the economics of digital markets. Economic literature has documented numerous economic bubbles across various markets and countries [50, 52, 42]. However, these markets are typically geared towards accredited investors, high-income individuals, or otherwise limited segments of the population, resulting in limited exposure for the general public. On the other hand, despite their complex technological underpinnings, the markets for NFTs and metaverse-related asssets are directly marketed to the ordinary investors—including children [60]. Faverio and Sidoti [26] shows that 17% of U.S. citizens, particularly young adults (e.g., 41% of young men), hold or have held cryptocurrencies. Given the similarities between cryptocurrencies and NFTs (common blockchain lineage, novel technology, etc.), we conjecture that cryptocurrency investors are also likely to bet on NFTs. Therefore, we argue that documenting the impact of the NFT and metaverse economic bubble could, more generally, help understand the effects of modern “digital bubbles,” which impact a segment of the population potentially broader than ever before.
To do so, we first examine relevant Reddit conversations to confirm that the Decentraland platform is representative of the metaverse and NFT ecosystem. Through topic modeling, we show that roughly 27% of the posts discussing Decentraland are about metaverse and NFT; the second-largest topic discussed relates to investments in metaverse projects (not limited to Decentraland). Even more interestingly, online activity (i.e., number of posts) related to these topics closely correlates with the trading volume of Decentraland land parcels. Furthermore, we observe that as land parcel prices are declining, violations of Reddit terms-of-use [^4] https://www.redditinc.com/policies/ in the relevant forums increase. This hints that users are getting frustrated by the economic downturn, and are forgoing civil discourse as the general interest from the broader online community is waning.
Second, we examine whether traditional (physical) real estate pricing models apply to Decentraland land prices to understand if at some time these prices reflected any traditional notions of the fundamental value of real estate. Decentraland uses a finite Euclidean topology to subdivide the entire world into parcels, each of which is paired to a unique NFT. Similar to holding a property title in the physical world, the owner of that NFT owns the parcel, and, accordingly, can modify its contents (e.g., build something), start a business, or sell it. Different from the physical world, though, Decentraland allows users to instantly teleport anywhere within the virtual world. The question then becomes whether parcel prices follow traditional real-estate price patterns, with prices decreasing as a function of distance from a desirable location. This would indicate a property characteristic of a “true” virtual city.
We show that in the early period when prices were low, parcel prices do in fact appear to follow this class of traditional real-estate models quite closely. But then, during the period of the rapid rise in land prices, location no longer appears to matter much–investors seem to purchase plots like they purchase other NFTs (e.g., art or collectibles), presumably for speculative purposes. In other words, land prices begin to deviate from the traditional real-estate models, suggesting a bubble-like phenomenon.
Our analysis of market demographics further shows that two distinct groups played major roles during the period we study. New entrants (i.e., those who started trading during the bubble), most of whom appear to be individual investors, were dominantly net-buyer groups (i.e., they bought more than they sold). On the other hand, early adopters provided liquidity (i.e., parcels) to the market and were mostly selling. This caused stark differences in monetary profits: Early adopters reaped 15,000 USD worth of profit per parcel on average. On the other hand, new entrants could rarely profit from the sales – in general, they lost in the order of 1,000 USD per parcel. In short, the time at which prices were high led to a disproportionate flow of money from new participants to early adopters.
Digital item marketplaces fostered by the metaverse and NFT are currently subject to lax or ill-defined regulations [25, 40, 56]. However, our results highlight significant financial risk to new market entrants, who may have been lured by the promise of riches, when the situation was markedly unstable. These new market entrants may not have had enough information to make rational or well-informed decisions. This result echoes the negative impact of the frenzy surrounding blockchain technology [17] and emphasizes the need for establishing appropriate standards of business conduct and improving user awareness, for example, through education or mandatory risk disclosure.
2 Related work and background
Economic bubbles
An economic bubble is broadly defined as a situation where a substantial rise in asset prices causes them to greatly exceed their fundamental value. Bubbles often follow inflated speculative trades and eventually result in significant investor losses [50, 33]. For cryptocurrencies specifically, Cheng et al. [17] shows that investments in blockchain companies during the 2017 initial coin offering (ICO) boom led to negative performance, and presumably to considerable investor losses. Due to their negative societal impact, the study of economic bubbles and of their root causes has been a focal interest of economics and monetary policy [5, 29]. A frequently cited cause of economic bubbles is that investors make biased decisions based on an extrapolation of past returns [15, 20, 21, 32]. As early as 1977, Minsky [46] claims that investors’ subjective extrapolation of past returns amplifies economic cycles (“financial instability hypothesis”). Aliber et al. [2] argue that economic bubbles evolve upon the arrival of good news leading speculative investors to overvalue asset prices. As a case in point, Cooper et al. [19] shows that, during the “dotcom” bubble of the 1990s, merely changing company names to include references to the Internet had a substantial positive effect on the corresponding stock prices.
Recent literature also proposes extrapolation-based models of economic bubbles [35, 8, 1, 30, 16, 23, 9]. For instance, Barberis et al. [9]’s model features a mixture of rational and extrapolating investors and describes the mechanism by which the latter dominate the market when good news is published, leading to an economic bubble. A remarkable feature of their model is that it explains the inflated trading volume during the bubble period.
Another thesis is that bubbles are caused by investors’ underestimation of downside risk. Gennaioli et al. [28]’s model show that unconscious exclusion of bad scenarios causes economic crises. Specifically, market participants ignoring bad scenarios leads to excessive issuance of risky assets. Unfortunately, as soon as market participants become aware of the downside risks, they abruptly withdraw funds, leading to steep declines in underlying asset value.
Bordalo et al. [14] propose “diagnostic expectation” as a mechanism to reconcile investors’ extrapolation and their negligence of downside risks. Their derivation stems from “representative heuristics” in the seminal work of Kahneman and Tversky [38]. Diagnostic expectation posits that investors weigh recent news too heavily in extrapolating from past returns, leading to overly optimistic (resp. pessimistic) forecasts after good (resp. bad) news. Maxted [43] incorporates diagnostic expectation to a real business cycle model and shows that optimistic (resp. pessimistic) outlooks cause increasing (resp. decreasing) asset prices, and that prices will revert to a rational level as optimism (resp. pessimism) fades away.
Diagnostic expectation may apply to the digital bubble we study. Indeed, media exposure of NFTs and of the metaverse abruptly surged in 2021, with notable events like the venerable Christie auction house holding an NFT auction, and Facebook’s rebranding itself as Meta [18, 45]. Investors may have expected a bright future for NFTs and for the metaverse, driven by the influx of news related to these concepts.
Metaverse and NFT
While the metaverse is now being featured as an extension of our current digital experience [41, 47], Linden Lab launched Second Life (SL) as an online virtual world that accommodates massive number of players as early as 2003 [13]. A feature that purportedly contributed to its popularity is the economy in which users could profit by trading their generated items or virtual real estate [57]. While Messinger et al. [44] suggest that making a profit is not the primary motivation for most users to play SL, Bakshy et al. [7] documented the existence of an active market in SL and found that sellers employ social ties in the virtual world to increase their revenue. As far as adverse effects, Stokes [59] points out the risk of money laundering facilitated by SL.
Second Life’s market structure resembles recent attempts to instantiate a Metaverse. On these platforms, users can profit by trading their created items and real estate published in the virtual worlds. A striking difference from earlier attempts, though, is that they employ blockchain as a core data layer where the state of the world, such as land or item custody, is maintained. For example, Decentraland has employed the Ethereum blockchain as a ledger for in-world items from its inception [51].
Public blockchains were originally introduced to implement a secure and transparent ledger system without relying on trusted third parties [49]. However, this technology is now supporting a large volume of mostly speculative cryptocurrency trading [11, 58, 39]. Besides cryptocurrencies, blockchains are also used to support non-fungible tokens (NFTs) [48]. Smart contracts, scripts to automate transactions, make secure and transparent tracking of each token’s ownership possible, by following established NFT standards [24]. In particular, even though all transactions are executed on publicly available ledger systems, NFT ownership is resistant to tampering, without relying on trusted third-parties. Hence, NFTs have been prominently touted as a technology that facilitates the trade of creative works in a secure and transparent manner. Another desirable feature of NFTs is that, being supported by decentralized blockchains, they are platform-independent. In particular, they are not tied to specific seller marketplaces: anybody can buy and sell them. As a result, metaverse platforms that use NFTs to represent in-world items allow these items to be traded outside the platform.
Our hope for this work is to complement the literature that has evidenced the possible negative impact of hype cycles surrounding blockchain technology [17, 37], by providing a quantitative analysis of the economic dynamics of the NFT market for metaverse-related items. More broadly, we hope to contribute to the pressing agenda of better understanding emerging “Web3.0” businesses [12].
3 Dataset
This section briefly describes Decentraland first and introduces our dataset. Our measurements and analysis are based on three data sources: (1) Ethereum blockchain records, (2) Decentraland public APIs, and (3) Reddit posts.
3.1 Decentraland dataset
Decentraland is an online platform developed since 2015. Public interest in Decentraland grew in 2017 when developers held an initial coin offering (ICO) of their MANA governance token along with the first land auction. In the auction, participants had the option to bid on parcels of land or stake their MANA toward community-initiated districts. [^5] See https://medium.com/decentraland/the-terraform-event-is-around-the-corner-bdda6a2d4367 and https://github.com/decentraland/districts. Following a second land auction in 2018 and the introduction of developer tools, the official Decentraland client was launched in February 2020. At the time of writing Decentraland has published roughly 93,000 parcels of land as an NFT collection called LAND.
Landowners can bundle up contiguous parcels they own into a new NFT token called Estate (EST), host their business or events on their parcels or estates, and list the parcels they own on (third-party) NFT marketplaces for sale.
Similar to real estate transactions, buyers can either buy parcels at the listed price, or bid a different price. They can also bid for unlisted parcels and directly negotiate with landowners.
Two different LAND NFT tokens only differ in the address (formally, a dimensionless integer) they represent in Decentraland. In particular, unlike many other NFT collections, there is no specific artistic value or artificial scarcity for a given LAND token. As such, we should be able to test whether real-estate pricing models and economic theories apply to these virtual parcels.
Ethereum blockchain records
Ether (ETH), the utility token for the Ethereum blockchain, has the second-largest market capitalization (after Bitcoin) at the time of writing. [^6] https://coinmarketcap.com/ Ethereum’s main feature is the ability to deploy smart contracts. In a nutshell, smart contracts are programs independently computed by all participants in the (Ethereum) network, who then collectively reach a consensus on the program output. This allows computations to be extremely robust to attacks. Decentraland employs smart contracts as a critical infrastructure. Because all Ethereum transactions, including all interactions with smart contracts, are recorded in a public ledger (the Ethereum blockchain), we can readily explore details of any Decentraland transaction, such as the contract functions that were called, the parties to the transaction, and any fees they might have incurred.
We collect all transactions interacting with the smart contracts listed on the Decentraland project page [^7] https://github.com/decentraland/contracts and third-party marketplaces where LAND and EST collections are listed until Dec. 31, 2023. While some user-created NFTs available on Decentraland, such as wearables, are hosted on the Polygon blockchain, all transactions on LAND and EST collections are hosted on Ethereum, [^8] https://docs.decentraland.org/player/market/marketplace/ so we focus on Ethereum. We validate our dataset by comparing it to Etherscan trade records. [^9] https://etherscan.io/token/0xf87e31492faf9a91b02ee0deaad50d51d56d5d4d\#tokenNFTTrade
Our dataset comprises four categories of transactions: (I) land purchases on the Decentraland marketplace, (II) land purchases on third-party marketplaces, (III) land auctions, and (IV) others (e.g., estate creation and dissolution). We store all information, such as the breakdown of sales prices per parcel, the creation/dissolution of estates, and ownership transfers for LAND token sales (category (I)) and for all transactions in category (IV). However, our data collection faces a limitation for EST token sales in category (I) and a subset of (II) and (III). Specifically, Decentraland held its first land auction off-chain, and at the time only relied on the Ethereum blockchain to collect deposits to participate in the auction, to refund deposits after the auction, and to transfer parcels to auction winners. Therefore, we cannot determine the exact price of each parcel if an auction participant won more than one parcel. Subsequent land auctions, bulk sales on third-party marketplaces, and estate sales share the same price ambiguity, i.e., we cannot determine the cost per parcel when somebody purchases more than one parcel. Bulk sales on third-party marketplaces may in particular include NFT items not related to Decentraland.
Decentraland APIs
Decentraland provides public APIs to collect parcel data, including parcel property information and ownership transfer history, sorted chronologically. [^10] https://github.com/decentraland/LAND-permissions-graph; https://github.com/decentraland/atlas-server. Every LAND token, including those bundled as EST tokens, has one or more property tags attached to it to denote the type of property the parcel refers to: 1) personally-owned lands can be bought and sold by users; 2) roads, and 3) plazas, on the other hand, are controlled by the Decentraland development team. (In particular, the Genesis plaza, where users spawn from, is in the center of the Decentraland map.) Finally, 4) districts denote a set of parcels developed by user communities based on public proposals. District founders have to obtain approval from the Decentraland DAO community to launch their district. There are currently 39 districts on the platform. A further 17 districts were dissolved in 2019.111111https://nftplazas.com/decentraland-districts/
Figure 1: Map of Decentraland color-coded with parcel types: district (grey), plaza (orange), ownable parcels (green), and roads (white).
Figure 1 shows the Decentraland map at the time of writing. We collect all the ownership transfer records, estate creation and dissolution records, and land property information until Dec. 31, 2023 from the APIs. Due to limitations on available property information, we cannot track back the addresses of districts dissolved in 2019.
Additionally, Decentraland provides a dataset containing the history of building construction on parcels. It has a comprehensive record of all buildings on parcels uploaded to the metaverse world. An entry for the dataset includes all the details about the construction, such as the uploaded date, its contents, and builder metadata.
3.2 Reddit dataset
Reddit is a major social media platform mainly used by young adults [6]. Reddit users can post in forums dedicated to specific topics (e.g., relationship advice or sports teams) called “subreddits.” For example, multiple subreddits focus on Decentraland, namely r/decentraland for general discussion, and r/MANACOIN for the MANA token. Figure 2 illustrates how users interact on Reddit. Discussions start on Reddit with a user submitting a post related to the subreddit’s theme. These posts, or “submissions” following standard terminology, include a title and a body of text. Other users can subsequently continue the discussion by responding to the submission. These responses only consist of a body of text (the title is that of the original submission) and are called “comments.” An important feature here is that administrators and content moderators for subreddits have the authority to remove posts (submissions and comments) that violate terms of use and/or rules for subreddits. Removal rules may vary across subreddits, but generally, inappropriate content such as scams, harassment, and more generally, posts considered abnormally rude are taken down from the platform.
We collect submissions and comments from 2019 through 2022 using the Pushshift Reddit dataset [10, 53]. In this dataset, removed posts are flagged as “removed” – while the text content is deleted, we still have access to metadata.
Figure 2: Reddit conversations. A user posts a “submission” under a relevant subreddit, and other users subsequently respond with “comments.” Content moderators and administrators monitor the content posted on subreddits.
3.3 Descriptive statistics
Decentraland
Table 1: Descriptive statistics for LAND and EST tokens.
2019 2020 2021 Q1 Q2 Q3 Q4 Q1 Q2 Q3 Q4 Q1 Q2 Q3 Q4 Number of transactions (LAND) Listing 7,616 6,522 2,813 2,928 7,629 3,386 2,323 2,241 3,361 3,211 1,695 4,164 Sales (Decentraland) 973 500 364 315 819 257 424 380 952 684 547 1,553 Sales (Third-party marketplace) 20 7 1 5 23 69 53 18 135 403 450 1,537 Bids placed 2,632 2,897 1,630 2,238 3,517 1,311 392 499 365 444 230 603 Bids accepted 96 195 93 99 330 173 53 56 51 65 8 52 Number of transactions (EST) Listing 3,594 3,002 1,379 1,454 3,248 1,608 1,120 921 1,161 1,189 755 1,504 Sales (Decentraland) 236 120 107 69 161 89 51 94 177 124 124 191 Sales (Third-party marketplace) 1 0 0 1 2 0 0 1 1 1 0 4 Bids placed 887 1,510 749 1,145 1,446 880 182 233 201 561 577 444 Bids accepted 23 86 48 36 68 56 24 23 20 59 51 56 LAND median price [$10^{3}$ USD] Listing 1.2 1.4 0.9 0.8 1.2 0.9 1.6 1.5 6.5 9.0 6.6 19.8 Sales (Decentraland) 0.6 0.7 0.5 0.4 0.8 0.6 0.8 0.7 4.0 6.0 5.0 15.4 Sales (Third-party marketplace) 0.7 0.5 0.6 0.4 0.5 0.5 0.8 0.5 4.5 4.3 4.2 14.4 Bids placed 0.7 0.7 0.5 0.4 0.6 0.5 0.8 0.7 3.2 5.5 4.2 10.0 Trading volume [$10^{6}$ USD] LAND – Sales (Decentraland) 0.8 0.5 0.3 0.2 0.8 0.2 0.4 0.4 4.7 4.6 3.1 26.5 LAND – Sales (Third-party marketplace) 0.0 0.0 0.0 0.0 0.0 0.1 0.1 0.0 0.7 3.2 3.8 24.2 EST – Sales (Decentraland) 0.9 0.6 0.3 0.2 0.6 0.4 0.2 0.3 4.1 4.5 3.2 17.8 EST – Sales (Third-party marketplace) 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 0.1 0.0 0.0 0.0 2022 2023 5-year Stats. Q1 Q2 Q3 Q4 Q1 Q2 Q3 Q4 Number of transactions (LAND) Listing 3,018 2,012 1,282 794 1,211 475 509 449 57,639 Sales (Decentraland) 1,010 477 214 212 167 81 88 93 10,110 Sales (Third-party marketplace) 1,197 724 364 395 335 221 190 284 6,431 Bids placed 458 559 402 345 125 100 75 44 18,866 Bids accepted 38 133 20 13 19 22 8 2 1,526 Number of transactions (EST) Listing 1,053 1,047 933 511 771 288 374 322 26,234 Sales (Decentraland) 155 109 74 67 35 23 37 35 2,078 Sales (Third-party marketplace) 22 31 40 47 6 0 6 5 168 Bids placed 1,112 411 1,131 1,486 241 305 276 124 13,901 Bids accepted 84 43 42 42 32 18 16 8 835 LAND median price [$10^{3}$ USD] Listing 17.3 8.9 5.1 2.9 3.0 2.0 1.3 1.4 2.0 Sales (Decentraland) 15.0 6.3 3.4 2.2 2.1 1.3 0.9 0.8 3.1 Sales (Third-party marketplace) 13.3 4.7 2.8 2.1 1.8 1.0 0.7 0.7 5.4 Bids placed 12.9 2.9 2.7 2.1 2.0 1.6 0.7 0.7 0.7 Trading volume [$10^{6}$ USD] LAND – Sales (Decentraland) 16.8 3.5 0.9 0.5 0.4 0.1 0.1 0.1 64.9 LAND – Sales (Third-party marketplace) 21.1 4.4 1.2 16.2 6.3 0.3 0.2 0.2 82.0 EST – Sales (Decentraland) 14.7 4.0 1.4 0.8 0.3 0.1 0.2 0.3 55.0 EST – Sales (Third-party marketplace) 0.0 0.1 0.0 0.0 0.0 0.0 0.0 0.0 0.5 * We exclude Decentraland sales whose price is higher than $10^{10}$ MANA as outliers.
Table 1 summarizes descriptive statistics about our parcel transaction dataset. From the number of transactions and trading volume, the number of sales transactions for LAND tokens is 5–8 times larger than that of EST tokens in late 2021, while their trading volumes are in the same order. Intuitively, this makes sense: EST prices are typically several times higher than LAND tokens since they hold multiple parcels. Buyers with smaller budgets, presumably individuals, are thus more likely to have purchased LAND tokens, while considerably fewer investors have enough purchasing power to acquire EST tokens.
Second, in the Decentraland LAND marketplace, the fraction of seller listings (row 1) that result in a sale (row 2) is higher than the fraction of (auction) bids (row 4) that result in successful purchase (row 5), in most of our measurement period. Intuitively again, desirable parcels are more likely to be sold at or close to seller asking price. On the other hand, if a parcel is only sold through the bidding process, it probably evidences a discrepancy between the seller’s expectations, and the buyers’ willingness-to-pay, making successful transactions rarer.
More generally, price distribution follows an expected order: listing prices are higher than direct sale prices, which are themselves higher than bidding prices. Among direct sales, those taking place on the Decentraland platform register higher prices than those on third-party marketplaces. This makes sense, as most desirable parcels generally immediately sell on Decentraland, while less popular land plots have to be shopped around more extensively, including on third-party websites. Direct sales registering for a higher price than accepted buyer bids is also unsurprising, as discussed above. Auctions typically are a “best offer,” whereas direct sales are much closer to the asking price.
Finally, starting in 2021, land prices show very large swings. The median price of a listing is in the order of USD 1,000 at the beginning of 2021, and jumped to close to USD 20,000 at the end of the year; actual sales prices went from USD 700 or so to about 15,000 USD. We note in particular a three-fold increase in sales price in Q4 only. However, the market quickly cooled down: by Q3 of 2022, sales prices had gone back down to about USD 3,000, and are back below USD 1,000 in 2023 Q4. In short, these prices movements seem to indicate that LAND tokens experienced a speculative bubble, which we will study further.
Reddit
Figure 3 shows the number of submissions on Reddit referring to “Decentraland” and the Google Trends score for the word “Decentraland” between Jan. 1, 2019 and Dec. 31, 2022. The number of submissions is highly correlated with Google Trends score (Pearson correlation coefficient = 0.75). This indicates that the Reddit dataset mirrors the general interest of web users in Decentraland.
Figure 3: LAND trading volume and number of Reddit posts. The blue bars and red line represent the number of submissions containing the word ”Decentraland” over two weeks and the Google Trend score for one week, respectively.
The substantial increase in Oct. 2021 is due to a bot-like submission recommending investment in Decentraland-related assets in a specific marketplace (see Appendix A). Therefore, we exclude these submissions from the analysis in subsequent sections.
4 Methods
To analyze the popularity and trading dynamics of Decentraland, we first explore topics discussed on Reddit. We then leverage blockchain records pertaining to Decentraland to carry out a regression analysis and compute profits and losses. Throughout this study, we use daily cryptocurrency price data from Yahoo! Finance to convert land prices denominated in cryptocurrencies to their corresponding USD value. [^12] https://finance.yahoo.com/crypto/.
Topic-modeling analysis of Reddit posts
We first investigate why Decentraland parcels gained popularity in 2021 from a topic-modeling analysis of Reddit posts. Specifically, we use the Gibbs-Sampling Dirichlet Multinomial Mixture (GSDMM) method [62]—a topic-modeling method frequently used for short text such as social media posts [54]—to identify topics of submissions referring to “Decentraland” either in the title or in the body. (See Appendix A for the details of GSDMM analysis.) For each topic identified by GSDMM, we count the number of submissions that match that topic. This allows us to identify what users are most interested in when discussing Decentraland, in particular, when the prices of Decentraland-related items (e.g., parcels) increases.
Starting in the second half of 2022, when prices start to fall significantly, an increasing fraction of comments are removed. This may give us clues about user frustration (as angrily venting may result in posts violating terms of service). Specifically, topic modeling allows us to better understand what users are most vociferous about.
Land price analysis
We next discuss our regression analysis on parcel prices. Our model is based on an economic theory that explains real-estate valuations, called bid-rent theory [3, 4, 27]. The fundamental idea is that the closer to a central business district (CBD), the higher the land price. This is because retail businesses view proximity to the CBD as more profitable than remote areas (due to the better access to customers), which in turn leads to increased competition between possible tenants and results in higher real estate valuation.
Bid-rent theory assumes the existence of a single center, which is the case in Decentraland. Users spawn from the center of the map (i.e., the Genesis Plaza in the center of Figure 1). So, businesses close to the Genesis Plaza should theoretically have more exposure to customers. A notable wrinkle compared to the physical world is that users can “teleport” to any parcel on the map once they spawn. In practice, though, we hypothesize that most users are primarily exploring the world by wandering around, mimicking the behavior of their real-world counterparts. On the other hand, parcel owners can develop their parcels very much like they would in the real-world: they can construct buildings, or host businesses such as NFT galleries [22] or casinos. In fact, our analysis of the number of visitors to parcels indicates that the hypothesis holds true for Decentralnad. See Appendix B for the details. Therefore, distance could still be an important metric in determining the land price both for sellers and buyers. (We will prove this is indeed the case in Section 6.)
We use this assumption to build an empirical regression, following Gupta et al. [31]:
$\log P_{i,n}=$ $c+\sum_{t}\left[\alpha_{t}+\delta_{t}\log\left(1+D_{i}\right)\right]Quarter_{t,n}$ (1)
$+\sum_{J}\beta_{J}I_{J,i}+\epsilon_{n}\ ,\ \ \epsilon_{n}\sim\mathcal{N}(0,\sigma^{2})\,,$
[…]
For our regression analysis, we will limit ourselves to listing and sales (i.e., purchase) prices for the LAND tokens on the Decentraland marketplace, and will not include third-party markets and EST token transactions in this analysis. We motivate this choice with four reasons: 1) sales on Decentraland dominate other categories of transactions (see Table 1); 2) transactions on Decentraland have no ambiguity in price breakdown, compared to those in third-party marketplaces that may include a LAND token as part of bulk offerings, possibly containing other, unrelated NFTs (see Sec. 3 for the details); 3) prior to 2021, third-party marketplace sales of LAND token were rare; 4) a number of bids appear to be unrealistic: they are several orders of magnitude lower than the median asking price, have therefore no chance to be accepted, and would only create noise in the regression.
Control variables
We have four control variables of interest. First, closeness to a street can be advantageous to bring in customer traffic to the parcel, thereby commandeering a higher land price. Moreover, in Decentraland, the profile page of a parcel prominently highlights information about its closeness to streets, districts, and Genesis Plaza, which may impact the parcel’s price. We capture these effects with three control variables:
$I_{Pub,i}=\left\{\begin{array}[]{ll}1&\mbox{if the closeness to $Pub$}\\ &\mbox{is displayed on}\\ &\mbox{the profile page of $i$-th parcel,}\\ 0&\mbox{otherwise.}\end{array}\right.$ (2)
where $Pub$ denotes one of {road, district, Plaza}.
We also notice that a limited number of parcels very close to the center of the map are exceptionally expensive compared to others and show a discrepancy in prices (see Appendix B). We account for this with a fourth control variable $I_{Close,i}$:
$I_{Close,i}=\left\{\begin{array}[]{ll}1&\mbox{if $D_{i}<20$,}\\ 0&\mbox{otherwise.}\end{array}\right.$ (3)
$I_{Close}=1$ for only 518 out of 44,480 ownable parcels (1.16%), so, for most parcels, any adjustment due to $I_{Close}$ will not be significant.
Profit and loss for land trading
Finally, we calculate the profit and loss for each unique Ethereum account involved in parcel trading. To do so, we will consider all LAND token sales transactions, including bids accepted in the Decentraland marketplace and sales in third-party marketplaces to track transfer records as completely as possible. We only exclude accounts that have transferred or received LAND tokens for free (to avoid reidentifying owners). For bulk sales, we estimate the price per parcel to simply be the total price divided by the number of NFT items involved. The inclusion of EST token transactions does not change results while a little noise is introduced.
5 Analysis of Reddit posts
Figure 4 shows word clouds for the three most common topics broached when discussing Decentraland – as categorized by the Gibbs-Sampling Dirichlet Multinomial Mixture (GSDMM) method. The most prominent topic is about the metaverse (not specific to Decentraland), NFTs, and cryptocurrencies; the second most prominent topic is investment advice regarding assets associated with metaverse projects; and the third most prominent topic concerns in-world events and features of Decentraland, such as fashion week, [^13] https://decentraland.org/blog/announcements/tradition-and-innovation-collide-decentraland-metaverse-fashion-week-2023 art week, [^14] https://decentraland.org/blog/announcements/metaverse-art-week-2023-the-metaverse-is-dead-long-live-the-metaverse and stores. These topics comprise 26.6%, 15.5%, and 13.6% of submissions, respectively. In short, Reddit users discussed first and foremost Decentraland in terms of a metaverse project leveraging blockchain technology, but also showed a secondary interest in profiting from the investment in metaverse-related items, and, to a lesser extent, an interest in features specific to Decentraland.
(a) First topic
(b) Second topic
(c) Third topic
Figure 4: The word cloud of the three most salient topics according to GSDMM. The size of words reflects their relative importance within the topic.
Figure 5 shows the temporal distribution of submissions categorized into the aforementioned topics. We observe interesting differences. The most salient topic, i.e., metaverse and NFT, increased in early 2021 and again in the 4th quarter of 2021, which were the times when NFTs gained substantial popularity, and Facebook rebranded itself as “Meta” in the hope that it would be the metaverse industry leader. In particular, the 4th quarter of 2021 features ten times more submissions than a year prior, indicating that Decentraland abruptly gained substantial popularity on Reddit, presumably due to the rising hype about NFTs and the metaverse. The second most popular topic (investment advice) shows an even more interesting distribution. In early 2021, the number of submissions categorized under this particular topic was the smallest among the main three topics of discussion. However, the number of investment-related posts experienced a significant increase and became the most actively discussed topic in the first quarter of 2022, denoting increasing excitement about investment opportunities. This coincides with the time when the metaverse gained renewed attention after Facebook’s announcement. A possible explanation is that that event sparked conversations about profiting from metaverse-, and by extension, Decentraland-related assets. In fact, the trading volume of Decentraland parcels is highly correlated with the number of submissions featuring these two most popular topics; the Pearson correlation coefficients are 0.74 and 0.83, respectively. On the other hand, submissions related to the third topic—in-world features—increased more moderately. Figure 5 shows that Decentraland features were more actively discussed than Decentraland-related investments before the end of 2021 – but the tide turned in late 2021, precisely as hype surrounding the metaverse started to build up. This suggests that this renewed interest in the metaverse primarily translated into investment-related (rather than feature-related) discussions, casting doubts user interest in actively participating in and building the platform.
Figure 5: Number of submission posts for the three most popular Decentraland topics under DMM analysis.
Figure 6 shows the number of top-level comments (i.e., replies to submissions) over time for the two most popular topics of discussion. As expected, the evolution over time of the number of comments tracks closely that of submissions (Figure 5). More interestingly, the ratio of removed comments increases sharply starting in Q1 2022 – both for Metaverse-related comments and for investment advice – reaching more than 60% in the first half of 2022. For comparison, the average removal rate on all of Reddit is 21.8% [34]. This surge in removed comments is temporally correlated with the number of transactions and prices of Decentraland land parcels plummeting (Table 1). While we do not have access to the text body of these removed comments, it seems highly likely that the sheer number of removed comments likely evidences user frustration about the poor performance of Decentraland assets.
(a) Metaverse/NFT (most popular topic)
(b) Investment advice (second most popular topic)
Figure 6: Comments on submissions categorized into the two main topics of discussion, and ratio of removed comments over time.
6 Analysis of Decentraland data
We next conduct the regression analysis of LAND token prices and then explore the profits and losses incurred by market participants.
6.1 Regression analysis
Figure 7 shows the spatial dependence of land prices ($\delta$) (see Appendix B for details). [^15] We excluded a small fraction of observations whose studentized residuals are more than 3.00 as outliers [61]. The grayed area corresponds to a 18-month window centered on 2021 Q4, which is when the hype bubble we identified earlier peaked. We first observe a clear difference between the “pre-bubble” situation from 2019 to early 2021 and the 2021–2022 bubble period. Specifically, before 2021, all prices show significant negative dependence – $\delta_{t}$ is around $-0.4$ or even lower for listing prices, indicating that sellers definitely account for location in their asking prices; buyers do, too, to a slightly lower, but nevertheless significant extent ($\delta_{t}\in[-0.2,-0.4]$). Concretely, parcels at a distance of 20 units from the Genesis Plaza ($\log D_{i}\simeq 3$) sell roughly for twice as much as the parcels at the edge of the map ($\log D_{i}\simeq 5)$.
However, Figure 7 shows a steep decline in the influence of geographical distance on purchase prices starting in 2021 Q2, when land prices per parcel increased to 5,000 USD, and NFT started to gain substantial popularity. $\delta_{t}$ even becomes statistically insignificant at the 5% level in 2021 Q4, meaning geographical distance has essentially no bearing on sales price at that time; this is when Decentraland abruptly gained increased attention as a metaverse platform. Interestingly, the trend somewhat reverses after the hype bubble has subsided (i.e., $\delta_{t}$ becomes statistically significant again, albeit less influential than pre-bubble). 2023 Q3 is an exception ($\delta_{t}\approx 0$), but this is probably due to the fact that the platform has basically become inactive: from Table 1, there were only 88 sales in that quarter, compared to several hundreds or more in earlier periods.
In an effort to check whether these findings generalize to other platforms, we analyzed the sales price of land parcels in The Sandbox, another major metaverse project leveraging blockchain technology. We found a similar decrease in the influence on prices of the distance from the center (see Appendix B for the details).
Figure 7: The spatial dependence of land prices ($\delta$) for quarters in 2019-2023. The error bars represent 95% confidence intervals (CIs). The grey-colored area shows the 18-month interval surrounding the “peak hype” of 2021 Q4.
6.2 Analysis of parcels’ value for commercial purpose
This section provides a deeper analysis of the inflating prices of Decentraland parcels in the bubble period. Decentraland parcels may have two distinct values. First, as discussed above, parcels have value as investment assets: the increasing prices could provide parcel owners with an opportunity for short-term capital gain. The other value lies in the potential for commercial activity in Decentraland. The surging popularity of Decentralnd would increase the number of active users, which is a significant chance for parcel owners to profit from commercial activities in Decentraland. We consider the contribution of the value for commercial activities with the building construction history on Decentraland parcels.
The primary commercial use of Decentraland parcels is advertising creative products, particularly artworks published as NFT items. Parcel owners place social media icons with links to their handles, where they can promote their creations in more detail. For example, Figure 8 shows a snapshot of a Decentraland parcel that has the links to its owner’s social media handles. When a player visits the parcel, they can hover over the icons and click on the links to visit the owner’s social media handles. We use the number of parcels that have links to social media profiles as an indicator of commercially used parcels and consider it as a measure of the prominence of commercial use of Decentraland parcels.
Figure 8: Snapshot of a Decentraland parcel with social media icons that have links to owner’s social media handles advertising their created items.
Figure 9 illustrates the number of uploaded scenes (i.e., the appearance of a parcel) that have social media links normalized by the number of parcel purchases and the number of developed parcels that have social media links normalized by personally owned parcels. [^16] The parcels do not belong to districts, Genesis Plaza, or roads. The normalization by parcel purchases helps us understand the proportion of parcel owners interested in promoting their creations compared to all owners. [^17] The number of purchases represents the minimum number of times parcels have changed ownership. As for the normalized number of developed parcels (lower figure), we assume that developed parcels are never cleared to vacant lots once a scene is uploaded, even if their owners change. Although this is a bold assumption, it aligns with rational behavior: while the owner may alter the scene, there is no rational incentive to clear the land and leave it vacant, as doing so does not increase the parcel’s value. This assumption implies that the number of developed parcels never decreases.
The figure depicts an interesting picture of commercial usage. To begin, the upper figure illustrates that only a small percentage of owners (5–10%) developed their holdings for advertisement. In particular, the ratio is only 4% in the fourth quarter of 2021. This suggests that they were not very interested in developing their properties to profit from commercial activities. The lower figure further emphasizes this point by showing that only less than 1.0% of properties were linked to their owners’ social media profiles in the fourth quarter of 2021. This indicates that commercial utilization of properties is not the primary driver of the price surge in late 2021.
Another important point illustrated in the figure is that the increase in the ratio does not align with the inflating prices of Decentraland parcels (Table 1). While parcel prices peaked in the fourth quarter of 2021 and have continuously decreased since then, the ratios shown in Figure 9 continue to increase until the second or third quarter of 2022. This indicates that the value of parcels for commercial use does not account for the change in parcel prices. Instead, the shift in the number of parcels linked to social media accounts suggests a change in owners’ motivation for purchasing parcels, from speculation in parcels to commercial purposes after the NFT and Metaverse frenzy subsides.
(a) The number of uploaded scenes ($N_{upload}$) normalized by the number of parcel purchases ($N_{Trade}$).
(b) The number of developed parcels that have social media links normalized by the number of personally owned ones. The blue bars show the number of developed parcels that have social media links ($N_{s}$) normalized by the number of personally owned ones ($N_{own}$). The red line shows the increase per quarter ($\Delta_{t}N_{s}/N_{own}$).
Figure 9: Building statistics on Decentraland parcels. Figure 9(a) shows the number of uploaded scenes that have social media links normalized by the number of parcel purchases. Figure 9(b) shows the number of developed parcels that have social media links normalized by the number of personally owned parcels.
Ruling out alternative hypotheses
We explored various alternative explanations for this result, beyond a lack of enthusiasm toward commercial opportunities in virtual worlds.
First, parcels could be valued for the social activities they could host, rather than their commercial potential—if so, customer exposure would become a much less important factor, and geographic dependencies would become marginal. However, the high parcel price during the booming period makes this alternative hypothesis highly unlikely: it is economically unreasonable that buyers would pay more than 15,000 USD for purely social activities (See Table 1 for median sales prices).
Another possible explanation is that the CBD of Decentraland expanded to the point of saturation, making the distance from the center less important. Namely, if all parcels were developed from nearby to distant locations, thereby making remote areas attractive to users just like areas near the center, the distance from the center would be less crucial. Our analysis rules out the possibility. We find that only 20-30% of parcels were developed by the end of the fourth quarter of 2021 (see Appendix B for the details). These data clearly reject the saturation hypothesis.
A third alternative interpretation is that many potential buyers were interested in commercial activity but were priced out of the market due to inflated parcel prices: the average parcel listing price tripled from 6,000 USD to 20,000 USD in a quarter (Table 1). However, there were still many wealthy enough buyers in the market. Specifically, Figure 10 shows three quantities related to buyers of multiple parcels. The bar plots show the fraction of transactions in which accounts that bought multiple parcels engaged (purchases in blue, sales in orange). We do not observe any meaningful decrease in purchases – 48.9% of sales in 2021 Q4 were to people buying more than one parcel. This suggests that some buyers have a large enough budget to buy parcels close to the center: the expected price of parcels close to the center is similar to that of remote parcels in 2021 Q4 as shown in Figure 7. Even more tellingly, the dashed line is the percentage of parcels bought by buyers of multiple parcels in a quarter and resold within two weeks or less (i.e., the ratio of short-term retention). We clearly observe an increase in “flipping” activity during the bubble: more than 40% of the parcels are sold within two weeks. In short, we can reject the hypothesis that buyers are priced out of the market. Instead, a large number of those wealthier, influential investors do not appear to invest in the long-term vision of a metaverse; instead, they merely attempt to make quick profits, which explains why valuations become disconnected from geographic considerations.
Figure 10: Sales and purchases of multiple-parcel buyers, and ratio of short-term ($<$ 2 weeks) holdings over parcels these buyers acquired in each quarter.
Another significant change in market demographics is the large inflow of new entrants. Figure 11 shows the cumulative number of unique sellers and buyers in quarters and its quarterly increase.
Figure 11: Cumulative number of unique sellers and buyers and its quarterly increase.
We see that there were 1,750 new buyers in 2021 Q4, three times the number for the previous quarter. Also, new entrants to that market in that quarter account for more than one-third of the cumulative number of unique buyers. Given our Reddit analysis, some of these buyers may be jumping on the metaverse bandwagon. We will see next whether this results in a profit.
6.3 Land sale profitability
Figure 12 shows landowner profit based on when LANDs token were purchased and sold. For instance, an account which buys LAND(s) in 2021 Q1 and sells it in 2021 Q2 is listed with a procurement date of 2021 Q1, and a sales date of 2021 Q2.
Figure 12: Profit landowners made based on when a parcel was bought (“procured”) or sold. The vertical black dashed line is the last week of Oct. 2021, when LAND price surged.
The figure shows a clear contrast between early adopters and the new entrants who joined during the bubble. Looking at profit as a function of the sales date, we observe that the largest profits ($\approx$ 10,000 USD) were for parcels sold in 2021 Q4 and 2022 Q1. After that, most sellers actually incurred losses in the order of 1,000 USD on average. When plotting profits as a function of the purchase date, we can see that early adopters who bought parcels at the beginning of 2021, followed by purchasers in 2020 and just before the price increase in late 2021, reaped the highest profits. Conversely, those who joined during the bubble (and especially right after the peak of October 2021) did not make significant profits—in fact, they barely broke even or even lost money.
Figure 13 further substantiates this finding. It shows the median profit per parcel and the number of unique sellers for parcels sold during 2021 Q4 and 2022 Q1 (upper) and in 2022 Q2 (lower).
Figure 13: Median sales profit per parcel, number of unique sellers, and number of parcels sold during selected quarters for the period of 2021 Q4 – 2022 Q1 (upper panel) and 2022 Q2 (lower panel). The light-blue area is the 15th-85th percentile.
Early adopters made significant profits during the bubble (upper panel); but for those who waited too long (lower panel) did not make as much. More importantly, people who joined after 2021 (i.e., new entrants possibly attracted by the metaverse hype bubble) and sold in 2022 Q2 lost money. The number of unique sellers (and sold parcels) also tells us that early adopters provided somewhat constant liquidity before and during the bubble; but once the bubble collapsed, that liquidity source dried up, and instead most sellers were new entrants that had joined during the bubble. The steep price decrease of 2022 Q1-Q2 led many of them to sell, driving prices further down, thereby perpetuating the cycle.
Figure 14 shows the number of unique landowners divided into tiers based on the number of parcels owned.
Figure 14: Unique landowners divided into tiers based on the number of parcels they own. The sum of the number of parcels held is always equal to the number of ownable parcels (i.e., 44,480).
More than half of the landowners (early adopters) held multiple parcels at the beginning of 2020. However, the number of single-parcel owners increased thereafter. In particular, we see an abrupt jump, of $\approx$1,000 single-parcel owners in 2021 Q4. On the other hand, the number of owners who possess more than five parcels of land does not change in that period. In other words, holders of large number of parcels sold land to single-parcel holders, most likely new entrants to the market.
In summary, the rapid appreciation of LAND tokens during the bubble profited early adopters who sold (some of) their possessions then. On the other hand, newly entrants who participated in land sales, probably motivated by the surge in the popularity of the metaverse, failed to reap profit from the sale and even frequently incurred losses. We see a transfer of wealth from new participants to early adopters—a situation reminiscent of what Soska et al. [58] had observed on BitMEX with larger portfolio holders continuously acquiring wealth from retail-level, less experienced investors.
7 Conclusion
This study has explored transactions on Decentraland, a decentralized virtual world platform. LAND tokens, which represent digital property title, saw their price skyrocket in 2021, which correlated with an increased amount of online discussions about Decentraland, specifically on Reddit. Topic modeling analysis revealed that discussions about Decentraland viewed that platform as representative of the NFT and metaverse ecosystems, and frequently focused on strategies to profit from these ecosystems. As a result, it is likely that Decentraland benefited from the hype surrounding NFTs and the metaverse. We performed regression and profit-and-loss analyses of land trades. We discovered that before the 2021 hype cycle, LAND prices followed established real-estate (bid-rent) theory, mimicking what we would see in the physical world. During that hype cycle, prices became disconnected from geographical considerations, leading the market to be subject to rampant short-term speculation—illustrated by the large amount of parcels being “flipped” over short periods of time, and characteristic of an economic bubble. Large landowners, with multiple parcels to their names, played an important role as both buyers and sellers at the same time. In the aftermath of the bubble, new entrants were left with unrealized financial losses reaching hundreds of dollars, or were forced to engage in fire sales to cut off their losses.
Our analysis shows that early adopters, who often had very large holdings, provided liquidity to the market and made sizable profits (10,000-15,000 USD per parcel, give or take) by selling some (but not all) of their holdings to new entrants that had arrived during the bubble. On the other hand, these new entrants made little or no profit – and even frequently lost money in the order of 1,000 USD.
To summarize, we can describe the 2021 Decentraland boom, likely fueled by the rising popularity of the topics related to NFT and metaverse, as a textbook case of bad investment by novice investors. Many new entrants tried to cash in on the hype, and failed. Did these new entrants have sufficient information to make sensible financial decisions? Platforms like Decentraland do not resemble traditional financial (or real-estate) investment platforms, which can lead to investor confusion in terms of not understanding the economic risks associated with them. As we demonstrated, these risks are real: how to mitigate them is an avenue for future work, that likely will include mandatory disclosures and user education.
8 Acknowledgements
This research was partially supported by Ripple’s University Blockchain Research Initiative (UBRI) at Carnegie Mellon and by the Carnegie Mellon CyLab Secure Blockchain Initiative. D.K. was supported by the Japanese Government Long-Term Overseas Fellowship Program. Some of the authors hold non-negligible cryptocurrency positions, but none on Decentraland. As part of exploratory efforts, some of the authors acquired a negligible amount of Decentraland real estate.
Appendix A Supplementary information about Reddit analysis
Bot-like message in Reddit submission posts
This section discusses bot-like submissions on Reddit.
Figure 3 shows a significant increase in the number of submissions in Oct. 2021. We address the cause of this increase in this section. Table 2 summarizes the top five words on the titles of Reddit submissions ranked by TF-IDF score, where TF-IDF scores for the $i$-th period are calculated as follows:
$V_{w,i}=TF_{w,i}\times\log\left(\frac{\left|\mathbf{D}\right|}{DF_{w}}\right)\,,$
where $TF_{w,i}$ and $DF_{w}$ are the term frequency of a word $w$ in the $i$-th observation period and document frequency of the word $w$, respectively. We normalize $V_{w,i}$ over words for each period: $\hat{V}_{w,i}\equiv V_{w,i}/\sqrt{\sum_{w}V_{w,i}^{2}}$.
Table 2: Top five words based on TF-IDF analysis for the titles of submissions with “Decentraland.” The values in parentheses represent the TF-IDF scores for words $\hat{V}_{w,i}$. The table on the left and that on the right summarize TF-IDF results for submissions referring to “Decentraland” and those excluding “coinsniper,” respectively.
TF-IDF for submissions w/ “Decentraland”
1st 2nd 3rd 4th 5th
’19 Q1-’20 Q4
virtual
game
ethereum
world
launch
’21 Q1
game
pizza
virtual
ethereum
atari
’21 Q2
poocoin
koing
pre
sale
ad
’21 Q3
ga
land
decentralize
property
sell
’21 Q4
coinsniper
lock
later
mc
lp
’22 Q1
metaverse
crypto
today
far
trend
’22 Q2
metaverse
bsc
tipsycoin
minute
project
’22 Q3
metaverse
crypto
today
trend
far
’22 Q4
crypto
today
trend
far
metaverse
TF-IDF for submissions w/ ”Decentraland” excl. those w/ “coinsniper”
1st 2nd 3rd 4th 5th
’19 Q1-’20 Q4
virtual
game
ethereum
world
launch
’21 Q1
game
pizza
virtual
atari
casino
’21 Q2
poocoin
koing
pre
sale
dynamic
’21 Q3
ga
land
decentralize
property
sell
’21 Q4
metaverse
minute
cryptocurrency
crypto
price
’22 Q1
metaverse
crypto
today
far
trend
’22 Q2
metaverse
bsc
tipsycoin
minute
project
’22 Q3
metaverse
crypto
today
trend
far
’22 Q4
crypto
today
trend
far
metaverse
The table on the left summarizes the results for submissions that refer to “Decentraland.” It shows that ”coinsniper” is ranked as the top result, followed by other seemingly unrelated words. We found that all the top five words in 4Q 2021 come from bot-like posts saying the same message. In fact, 1,350 out of 1,788 posts (75.5%) submitted in the week of Oct. 24, 2021, are a single message that recommends investment in Decentraland-related cryptocurrencies and NFT collections in CoinSniper, a cryptocurrency investment platform featuring newly emergent cryptocurrencies, reflecting the surging price of MANA token and parcels in Decentraland. Moreover, 1,989 posts in that quarter had the same message, which indicates bot activity. The table on the right summarizes the TF-IDF result for the submissions referring to “Decentraland,’ excluding those referencing “coinsniper” too. It appears that “metaverse” has the highest rank, followed by other relevant words. This seems to be a reasonable result. Therefore, we omit the submissions referring to “coinsniper” from GSDMM analysis.
Details of DMM analysis
[…]
Appendix B Supplementary information about regression analysis
This section describes supplementary information on our regression analysis.
Distance dependence of the number of visitors to parcels
This appendix considers the dependence of the number of visitors to parcels on the distance from the center of Decentraland map. Decentraland provides the number of visitors to parcels over past 30 days. We collect the data for the period from Apr. 15, 2024 to June 5, 2024.
Figure 15 shows the median number of parcel visitors. To calculate the median numbers, we first find the average number of visitors for each parcel over time. Then, we calculate the median number of visitors for the parcels grouped by distance in 10-unit blocks. The result shows a decrease in the number of visitors as the distance from the center increases. While we cannot directly assess whether this negative relationship holds true for the period from 2020 to 2023, this result suggests that the number of visitors negatively depends on the distance from the center as we hypothesize in Section 4.
Figure 15: The median number of parcel visitors for the period from Apr. 15, 2024 to June 5, 2024. To calculate the median numbers, we first find the average number of visitors for each parcel over time. Then, we calculate the median number of visitors for the parcels grouped by distance in 10-unit blocks. The area colored light blue represents the 25th to 75th percentile range for the blocks.
Price dependence on the distance in the close region
Figure 16 shows the observations of listing transactions on the Decentraland in 2022.
Figure 16: The observations of listing transactions on the Decentraland marketplace in 2022. The blue points and red points show individual listing transactions and the average price for every 0.25 intervals in log distance, respectively.
Prices for parcels close to the center ($\log(1+D)<3\Leftrightarrow D<20$) are more expensive than those in remote areas, so there appears to be a price discrepancy. Both listing and sales prices demonstrate similar trends. Hence, we have to consider a control variable ($I_{close}$) to mitigate the influence of the discrepancy.
Detailed regression result for land parcel prices
This paragraph summarizes the supplementary information about our regression analysis. We apply Eqn. (2) for each one-year period between 2019 and 2023 to accurately identify outliers in the process described below. In the regression analysis, we found that a small fraction of observations have unrealistically high or low prices, typically several orders of magnitude different from median sales/listing prices. To exclude such noisy observations, we excluded the observations whose studentized residuals are more than 3.00 as outliers [61].
Table 3 summarizes the result of our regression analysis. We perform estimations using the ordinary least square method (OLS).
Generalizability of our results
A critical question to our regression result is whether the decrease in the negative dependence on the distance in 2021 and 2022 is generalizable to other platforms. To look into the question, we consider The Sandbox, another major metaverse platform that adopts NFT as a core foundation of land parcels, publishing them as an NFT collection. A stark difference between Decentraland and The Sandbox, though, is that the parcels in The Sandbox are gateways to separate metaverse worlds. Instead of a single “flat” metaverse world in Decentraland, users move their mice over parcels on the map to go to metaverse worlds whose links show up once they click parcels in The Sandbox. This mechanism is different enough from physical real estate that it is unclear whether real-estate pricing theories, such as bid-rent theory, are relevant.
However, even with the limitation, parcels close to the center of The Sandbox are advantageous in customer exposure since they are displayed when users visit The Sandbox. Hence, we measure how sales prices of The Sandbox’s land parcels are affected by their distance from the center. In our analysis, we focus on transactions performed using the “atomicMatch_” function in OpenSea’s contract. [^19] https://etherscan.io/address/0x7Be8076f4EA4A4AD08075C2508e481d6C946D12b This is because the majority of sales in 2021 and 2022 were carried out on OpenSea, and 89.4% of them used the function. Due to the migration of The Sandbox from Ethereum to the Polygon blockchain, [^20] https://www.sandbox.game/en/blog/the-sandbox-is-deploying-on-polygon/3060/ which was announced in June 2022, the data is limited to the period of Q1 2020 to Q3 2022. We use a simple regression formula,
$\log P_{i,n}=c+\sum_{t}\left[\alpha_{t}+\delta_{t}\log\left(1+D_{i}\right)\right]Quarter_{t,n}\,,$ (4)
where we take the geographical coordinate data from The Sandbox’s publicly available API. The definition of terms in Eqn. (4) is the same as in Eqn. (1). Figure 17 shows the dependence of sales prices ($\log P$) on the distance from the center. It shows that geographical distance matters even less than outside of the price bubble in Decentraland. This result hints that our finding for Decentraland can be applied to other metaverse platforms as well.
Figure 17: The spatial dependence of land prices ($\delta$) for quarters in 2020-2022. The error bars represent 95% confidence intervals (CIs). The grey-colored area shows the 18-month interval surrounding the “peak hype” of 2021 Q4.
Discussion about saturation hypothesis
This appendix examines the spatial distribution of developed parcels to determine whether the saturation hypothesis accounts for the surge in prices in late 2021. An important consideration in distribution analysis is the relationship between the number of parcels and distance. It is crucial to note that the number of parcels per distance increases proportionally with distance. Comparing the number of developed parcels without considering this factor will lead to an overestimation of development in remote areas. Therefore, we normalize the number of developed parcels within a distance by the total number of parcels in that distance.
Figure 18 shows the number of developed parcels ($N_{Dev}$) normalized by the total number of parcels ($N_{Tot}$) for every 10-distance interval over half-year periods from 2020 to 2023.212121Due to the assumption that developed parcels are never cleared (see Sec. 6.2), the number of developed parcels increases monotonically over time. It shows the maximum number of parcels developed by the time It shows that only 20-30% of parcels in the map were developed by 2021 Q4, suggesting that the CBD of Decentraland is far from saturation.
Figure 18: The spatial distribution of developed parcels. Lines show the number of developed parcels ($N_{Dev}$) normalized by the total number of parcels ($N_{Tot}$) for every 10-distance interval over half-year periods from 2020 to 2023.
Table 3: Regression results. The values in parentheses represent the standard errors for estimates.
2019 2020 2021 2022 2023 Q1 Q2 Q3 Q4 Q1 Q2 Q3 Q4 Q1 Q2 Q3 Q4 Q1 Q2 Q3 Q4 Q1 Q2 Q3 Q4 Listing Price c 8.767 (0.090) 8.676 (0.086) 10.226 (0.130) 11.378 (0.134) 9.036 (0.198) $\alpha_{t}$ 0.083 (0.122) -0.906 (0.159) 0.493 (0.163) -0.672 (0.137) 0.779 (0.146) 0.371 (0.157) -0.591 (0.174) -0.189 (0.217) 0.704 (0.169) -1.015 (0.182) -1.725 (0.215) -2.518 (0.256) -0.992 (0.345) -1.375 (0.346) -0.269 (0.387) $\delta_{t}$0.449 (0.019) -0.430 (0.020) -0.298 (0.031) -0.423 (0.031) -0.385 (0.018) -0.302 (0.026) -0.516 (0.028) -0.436 (0.030) -0.388 (0.029) -0.166 (0.029) -0.310 (0.040) -0.257 (0.026) -0.359 (0.029) -0.310 (0.032) -0.249 (0.042) -0.188 (0.051) -0.283 (0.043) -0.142 (0.065) -0.133 (0.068) -0.352 (0.076) $I_{District}$ 0.160 (0.013) 0.222 (0.012) 0.178 (0.017) 0.143 (0.021) 0.153 (0.033) $I_{Plaza}$ 0.311 (0.017) 0.247 (0.017) 0.208 (0.024) 0.105 (0.028) -0.030 (0.045) $I_{Road}$ 0.356 (0.019) 0.178 (0.018) 0.221 (0.023) 0.188 (0.028) 0.451 (0.044) $I_{Close}$ 0.375 (0.049) 0.747 (0.063) 1.066 (0.065) 0.899 (0.067) 1.522 (0.128) Adjusted $R^{2}$ 0.181 0.167 0.318 0.440 0.274 Num. of obs. 19,876 15,579 12,073 6,288 2,132 Log-likelihood -24,379 -17,468 -15,716 -7,239 -2,321 AIC $4.878\times 10^{4}$ $3.496\times 10^{4}$ $3.146\times 10^{4}$ $1.450\times 10^{4}$ $4.67\times 10^{3}$ BIC $4.89\times 10^{4}$ $3.51\times 10^{4}$ $3.15\times 10^{4}$ $1.46\times 10^{4}$ $4.74\times 10^{3}$ Prob(F-stat.) 0.00 0.00 0.00 0.00 0.00 Sales Price c 7.925 (0.174) 7.814 (0.172) 10.027 (0.151) 9.976 (0.113) 8.347 (0.328) $\alpha_{t}$0.861 (0.274) 0.146 (0.292) -0.966 (0.302) -0.411 (0.334) 0.286 (0.245) 0.178 (0.320) -1.060 (0.223) -0.852 (0.257) -0.424 (0.188) -0.503 (0.198) -0.434 (0.296) -1.795 (0.261) -0.067 (0.503) -1.615 (0.571) -0.671 (0.510) $\delta_{t}$0.363 (0.037) -0.165 (0.047) -0.428 (0.054) -0.237 (0.055) -0.260 (0.036) -0.239 (0.062) -0.307 (0.041) -0.293 (0.058) -0.431 (0.033) -0.097 (0.038) -0.170 (0.046) -0.029 (0.026) -0.103 (0.024) -0.182 (0.037) -0.287 (0.061) -0.111 (0.053) -0.182 (0.069) -0.231 (0.086) 0.016 (0.109) -0.204 (0.042) $I_{District}$ 0.103 (0.023) 0.163 (0.023) 0.117 (0.019) 0.046 (0.018) -0.034 (0.042) $I_{Plaza}$ 0.136 (0.038) 0.129 (0.037) 0.105 (0.027) 0.152 (0.026) 0.118 (0.060) $I_{Road}$ 0.280 (0.028) 0.029 (0.028) 0.115 (0.022) 0.151 (0.023) 0.186 (0.055) $I_{Close}$ 0.697 (0.107) 0.725 (0.179) 0.635 (0.104) 0.270 (0.110) 1.332 (0.277) Adjusted $R^{2}$ 0.267 0.169 0.571 0.770 0.520 Num. of obs. 2,151 1,877 3,708 1,712 305 Log-likelihood -1,567 -1,265 -2,838 -602 -103 AIC 3,157 2,554 5,699 1,228 230 BIC 3,225 2,621 5,774 1,293 275 Prob(F-stat.) 0.00 0.00 0.00 0.00 0.00